Please help me to solve this system algebraically:
1/2x+2/3y=5/6
5/12x+7/12y=3/4

Answer by mananth(16946)

1/2 x+2/3y=5/6
LCD is 6
3x+2y /6 =5/6
3x+2y=5*6/6
3x+2y=5...................1
5/12 x+7/12 y=3/4
5x+7y /12= 3/4
5x+7y= 12*3/4
5x+7y=9...................2
..
Multiply equation 1 by -5 and equation 2 by 3 and add them
-15x-10y+15x+21y=-25+27
11y=2
y = 2/11
plug the value of y in equation 1
3x+2*2/11 =5
3x=5-4/11
3x=51/11
x= 17/11